OpenStudy (anonymous):

Write an expression that best represents the context of the word problem. Solve for the unknown values. If Sarah is 24 years younger than her mother and if the sum of their ages is 68, how old is Sarah?

OpenStudy (raden):

the expression algbera of this, is s = m + 24 and s + m = 68 with s = sarah's old, m=mother

OpenStudy (anonymous):

So, the answer would be?

OpenStudy (raden):

sorry, the first should s=m-24 then to get s, do subtituting or elimination method from s=m-24 ---> m=s+24 substitute m=s+24 to s+m=68, so s+m=68 s + s+24=68 2s + 24 = 68 2s = 68-24 2s = 44 s = 44/2 = 22